# REGULAR ARTICLE

### v1

@@ 3.0.2 Jarzynski equality (1997):
$F(\beta,\lambda)=-(1/\beta)\ln{\sum_{\Gamma\in D_{\lambda}}e^{-\beta H(\Gamma,\lambda)}}.$ (30)
@@
Since, for a regular protocol, the dynamical image of $D_{0}$ after time $\tau$ becomes $D_{1}$, the normalisation of $p_{f}(\Gamma_{1})$ is guaranteed. So, we can write [27]
@@
This remarkable property, which applies for the distribution of work and several other thermodynamic quantities, is known as the integral fluctuation theorem (IFT) [24, 27]. From Eqs. (32)-(34), we identify
@@
with $\Delta F=F(\beta,\lambda(\tau))-F(\beta,\lambda(0))$. Substituting Eq. (35) into Eq. (34), the corresponding IFT reduces to the celebrated Jarzynski equality (JE) [33]:
@@
This equality relates an average over the work required in this process to the free-energy difference between the two canonical equilibrium states: one corresponding to the value $\lambda(\tau)$ of the control parameter at time $\tau$ and other to its initial value $\lambda(0)$, both taken w.r.t. the same inverse temperature $\beta$. Further, applying Jensen’s inequality to Eq. (36), gives $\mathcal{W}=\langle W(\Gamma_{0})\rangle\geq\Delta F$ [24]. If we define the dissipative work as $W_{\text{diss}}=W-\Delta F$, this inequality can be rewritten as $\langle W_{\text{diss}}\rangle\geq\ 0$ [27]. For smaller systems, unlike in the thermodynamic limit, $W_{\text{diss}}$ need not always be positive for individual realisations of the process; it may even be negative. Nevertheless, upon averaging, it becomes consistent with classical thermodynamics. If we identify $W_{\text{diss}}=T\Sigma$, where $\Sigma$ is the entropy production, then combining the IFT with Jensen’s inequality leads to $\langle\Sigma(\Gamma)\rangle\geq 0$ which is recognised as the second law of stochastic thermodynamics [25]. The application of JE to mesoscopic systems in modern times has been remarkable. Notably, JE places no restriction on the speed of the transformation and thus remains valid even far from equilibrium. This makes it a powerful tool for determining free-energy differences in molecular dynamics simulations, driven harmonic oscillators, and even real biomolecular experiments [28]. Many researchers view JE as having brought nineteenth-century thermodynamics into the twenty-first century, given its profound implications [21].
3.0.3 Crooks Theorem (1999):
@@
$\begin{split}\mathcal{P}_{R}(-W)=&\sum_{\Gamma^{R}_{0}\in D_{R}}\delta(W_{R}(\Gamma^{R}_{0})+W)e^{-\beta[H(\Gamma^{R}_{0},\lambda(\tau))-F(\beta,\lambda(\tau))]}\\ =&\sum_{\Gamma_{0}\in D_{F}}\delta(-W_{F}(\Gamma_{0})+W)e^{-\beta[H(\Gamma_{0},\lambda(0))+W_{F}(\Gamma_{0})-F(\beta,\lambda(\tau))]}\\ =&\mathcal{P}_{F}(W)e^{-\beta(W-\Delta F)}\end{split}$ (37)
Finally, we can write [35, 36]
@@
The CFT relation in Eq. (38) is stronger than JE [cf. Eq. (36)], since JE can be derived by integrating both sides of Eq. (38) and using $\int dW\mathcal{P}_{R}(-W)=1$. Collin et al. (2005) [37] experimentally verified Crooks’ theorem in a single-molecule RNA hairpin folding–unfolding experiment using optical tweezers. In this experiment, $W_{\text{diss}}$ was sometimes negative, but on average it was positive. Thus, for small biological systems, this theorem provides a clear-cut framework for calculating free-energy changes from work distributions.
3.0.4 The different DFT and furthermore:
@@
which is an example of the detailed fluctuation theorem (DFT) [27]. The DFT is a stronger result than the IFT. While the IFT provides a single global constraint on the work distribution (and more generally on entropy production), the DFT establishes a local balance relation between the probabilities of observing a trajectory and its corresponding antitrajectory. Consequently, in a steady-state system, it is sufficient to know $\mathcal{P}(W)$ for positive values of $W$ in order to reconstruct the entire distribution [25]. Both the JE and the CFT can be regarded as special cases of the IFT and DFT when the initial and final states are equilibrium states.
@@
There exists a vast literature on FTs, depending on the observables of interest, the underlying dynamics of the system, and the hypotheses or models considered. Evans and Searles [30] first proposed such a relation. Subsequently, Gallavotti and Cohen [38] derived a detailed fluctuation theorem for the entropy production rate in deterministic chaotic systems. Later, the JE [33] and CFT [36] were developed for systems driven between two equilibrium states, becoming landmark results that enabled equilibrium free-energy differences to be estimated from non-equilibrium measurements. This also led to the rediscovery of earlier works by Bochkov and Kuzovlev [39], although their definition of work differs from the present one.
@@
Moreover, introduction of trajectory probabilities for forward and backward processes naturally gave rise to a variety of FTs. For entropy production, if we separate its contributions into two parts — one arising from time dependence and another from external driving (adiabatic and non-adiabatic entropy production) — the IFT leads to the Hatano-Sasa theorem [40], which constitutes a cornerstone in the generalisation of FDR (fluctuation-dissipation relation) to non-equilibrium steady state. Subsequently, Rao and Esposito [41] developed a unified framework that systematically derives many of the FTs reported in the literature. In addition to work and entropy, considerable attention has been devoted to the fluctuation symmetry of heat [42, 43]. Jarzynski and Wójcik [44] analysed the statistics of heat exchange between two systems coupled to distinct thermal baths, deriving a fluctuation relation that is now widely referred as exchange fluctuation theorem (XFT/EFT) [45]. More recently, the study of joint statistics of thermodynamic quantities and their associated FTs has attracted considerable attention. Interestingly, while the individual quantities may or may not obey an FT on their own, their joint probability distribution often satisfies a joint fluctuation theorem, thereby revealing richer symmetry structures in far from equilibrium thermodynamics [46, 47].
@@
So far, we have provided an account of the development of FTs in a classical settings. To understand their quantum counterparts, however, we must dig deeper. The key difference arises because any measurement of quantum states or observables disturbs the system and its trajectories, unlike in classical systems. In fact, continuous monitoring can even freeze the dynamics due to the quantum Zeno effect [48]. In the quantum regime, the very definitions of work, heat, and entropy must be handled with care. Thus, exploring quantum versions of FTs is a rich and promising area of current research.
4 Linear Response Theory and Quantum Work Fluctuations
@@
— James Clerk Maxwell [49]
Up to this point, we have discussed the role of fluctuations in a thermodynamic setup, where general thermodynamic quantities vary due to interactions with heat bath. In 1957, Ryogo Kubo [50] laid a concrete foundation within the Hamiltonian framework, showing how any arbitrary dynamical observable of a system coupled to a thermal bath deviates from equilibrium under an external drive, and how irreversible macroscopic behavior appears as a direct consequence of microscopic fluctuations [51] when observed through the lens of linear response of the system. Since then, Linear Response Theory (LRT) [52, 53, 54, 55] has held a central place as a powerful tool for studying near-equilibrium systems, with far-reaching applications in fluid dynamics, transport phenomena (both classical and quantum), condensed matter physics, many-body problems, chemical kinetics, and beyond [56]. More recently, it started elucidating its impact on quantum thermodynamics (QTD) [57, 20].
4.0.1 Introduction to LRT:
Here, we establish the Kubo formulas in parallel for both classical [55, 58, 52, 54, 53] and quantum mechanics [52, 54]. These formulas describe correlation of fluctuations, such as response and relaxation functions between observables of different dynamical variables. We discuss the properties and applications of these response and relaxation functions in the context of QTD.
@@ 4.0.1 Introduction to LRT:
We now perturb the system, starting from an initial time $t_{0}$ (eventually we take $t_{0}\to$-$\infty$) by applying a time-dependent external force $\lambda(t)$. This force is spatially homogeneous and couples with an arbitrary observable $A$ of the system. The perturbed Hamiltonian is then $H(t)=H_{0}+H^{\prime}(t)=H_{0}-\lambda(t)A$. For $t\geq t_{0}$, the density operator (or the phase-space distribution in the classical case), $\rho(t)$ evolves according to the Liouville–von Neumann equation [59, 56]
@@
Here, $\{\,\cdot\,$$\cdot\,\}$ and $\cdot\,$$\cdot\,$] denote the Poisson bracket and the commutator, respectively [56]. This formulation allows us to treat classical and quantum dynamics in a unified manner. With this choice one recovers the standard forms:
@@
Since the Liouvillian is Hermitian (one can show $\mathcal{L}=\mathcal{L}^{\dagger}$), the operator $e^{\pm i\mathcal{L}t}$ serves as the time evolution operator [52]. We now investigate the evolution of the density operator under an external perturbation in the linear regime. The Liouvillian can be split as $\mathcal{L}=\mathcal{L}_{0}+\mathcal{L}_{\text{pert}}$, associated with $H_{0}$ and $H^{\prime}$ respectively. The perturbed density operator is written as $\rho(t)=\rho_{\text{eq}}+\delta\rho(t)$. Inserting these into Eq. (43), we obtain [55, 53]
@@
where, the response function is defined as [50, 52]
@@
The relaxation function is then defined as [50, 52]
@@
which characterizes how $B$ relaxes after the perturbation is removed. Using Kubo’s identity for the canonical ensemble [56, 60],
@@
This function is also referred to as the canonical correlation function. However, in quantum mechanics, it is often more useful to express the relaxation function in a form that accounts for the non-commutativity of operators, leading to the symmetrised correlation function, denoted as $\Psi_{AB}(t-t^{\prime})$ [52]. Since it accurately captures the response (i.e., the fluctuation from equilibrium) caused by an external disturbance through time correlations between two dynamical observables, LRT has become increasingly important in quantum thermodynamics, where the control [61] and optimisation of quantum devices governed by fluctuations [62], particularly quantum fluctuations, are of growing interest. In our context, the autocorrelation function ($A=B$), which measures how an observable correlates with itself at different times, is of central importance and will be discussed in the next section.
4.0.2 Quanum Work Fluctuations and LRT:
@@
We live in an era where we aim to reach, utilise, and design systems on scales that have never been achieved before [63, 64, 65, 66]. To fully understand and optimise machines at the quantum level, where thermal and quantum fluctuations play equally important roles, we need to deepen our theoretical understanding of FTs, a.k.a., quantum fluctuation theorems (QFTs) [67, 68, 69, 70]. While QFTs can be formulated for various thermodynamic quantities, in this article we focus exclusively on work. Within the LRT framework, formulating QFTs for dissipative work has yielded new insights [71]. More recently, Guarnieri et al. [57] demonstrated that even near equilibrium, the work distribution is strongly influenced by quantum fluctuations, which are captured through the relaxation functions.
@@
For example, consider a quantum system with Hamiltonian $H_{0}$, initially thermalised in the normalised Gibbs state $\rho_{\text{eq}}=e^{-\beta H_{0}}$. The system is then perturbed unitarily as $H_{t}=H_{0}-\lambda(t)A$ over $t\in[0,\tau]$, where $A$ is an observable and $\lambda(t)$ is a weak dimensionless driving protocol ($\lambda(t)\ll 1$, $\|A\|=1$). At $t=\tau$, when the drive is switched off, the relaxation of $A(\tau)$ is captured by the symmetrized autocorrelation function $\Psi_{0}(=\Psi_{AA})(t-t^{\prime})$, which encodes correlations of $A$ at different times. Since changes in $A$ also modify the Hamiltonian, $\Psi_{0}$ links correlations between $H_{\tau}$ and $H_{0}$, providing a natural connection to the very definition of quantum work [57, 72].
@@
The most accomplished and widely used definition of work in QTD is the two-time projective measurement [69, 21] (TPM) scheme, where the system’s energy is measured at the beginning and end of the protocol. At $t=0$, a projective energy measurement is performed, then the system evolves under the time-dependent Hamiltonian, and at $t=\tau$ a second measurement is made. In this framework, work $W$ is a stochastic variable. For a single realization, it is given by $W[|m\rangle;|n\rangle]=E_{n}(\lambda(\tau))-E_{m}(\lambda(0))$, where $|m\rangle$ and $|n\rangle$ are the initial and final eigenstates with energies $E_{m}(\lambda(0))$ and $E_{n}(\lambda(\tau))$, respectively. The probability distribution of work is obtained by averaging over an ensemble of realizations [73]
@@
where $p(|m\rangle\rightarrow|n\rangle)$ denotes the probability of observing the transition $|m\rangle\rightarrow|n\rangle$. If the system is initially in a thermal Gibbs state $\rho_{eq}$, Eq. (55) reduces, as shown by Scandi et al. [74], to the standard TPM definition of work:
$\mathcal{P}(W)=\sum_{W=E_{n}-E_{m}}\langle m(\lambda(0))|\rho_{eq}|m(\lambda(0))\rangle\,\big|\langle m(\lambda(0))|n(\lambda(\tau))\rangle\big|^{2}.$ (56)
This expression represents the product of the probability of the system initially being in state $|m\rangle$ with the transition probability $|m\rangle\rightarrow|n\rangle$. However, since each projective energy measurement perturbs the system — and may even destroy its state — the full $\mathcal{P}(W)$ cannot be directly accessed in practice through this scheme [74, 21].
However, from the statistical theory, we know that any probability distribution can be reconstructed from its cumulant generating function (CGF) [69], defined as
$K(\lambda)=\ln{\langle e^{-\lambda\beta W}\rangle}:=\ln{\int_{-\infty}^{\infty}}dW\mathcal{P}(W)e^{-\beta\lambda W},$ (57)
where $\lambda=\lambda(\tau)$, denotes the final value attained by the protocol. The CGF is additive under independent random measurements, and allows one to recover the probability distribution $\mathcal{P}(W)$ through an appropriate inverse Fourier transform. Moreover, all the statistical information of $\mathcal{P}(W)$, such as cumulants, can be computed by differentiating the CGF [75]:
@@
Since $W=\Delta F+W_{\text{diss}}$, the CGF naturally separates into two parts: a deterministic term depending only on the free-energy difference, and dissipative contribution due to the irreversible process, as shown by Scandi et al. [74]
@@
where $K^{\text{diss}}(\lambda)=\ln{\int_{-\infty}^{\infty}}dW_{\text{diss}}\mathcal{P}(W_{\text{diss}})e^{-\beta\lambda W_{\text{diss}}}$. So, the most interesting quantity in QTD is the dissipated work, $W_{\text{diss}}$, which quantifies the irreversible cost of driving the system out of equilibrium. Its cumulants encode the full statistical information relevant to QFTs.
Guarnieri et. al. [57], building on the elegant method of Shitara and Ueda [76], established a landmark bridge between LRT and QFTs as
@@
where a function of $\lambda$, $g(\lambda)$, has to be convoluted with our old friend $\Psi_{0}$ to get the CGF of the dissipated work. The symmetry of $g(\lambda)=g(1-\lambda)$ ensures $K^{\text{diss}}(\lambda)=K^{\text{diss}}(1-\lambda)$ [57]. Operating an inverse Laplace transform on it, gives our pre-acknowledged companion
@@
a special form of ESFT [57], now established in the quantum regime. The essential finding of Eq. (60) is that any phenomenological relaxation function can fully determine quantum work distribution in the linear regime. Moreover, this framework predicts all higher-order cumulants are positive, i.e., $\kappa_{W}^{(n)}\geq 0$ [57], leading to non-Gaussian tail (such as skewness and excess kurtosis), driven by quantum fluctuations [74]. To quantify these non-classical effects, Guarnieri et al. [57] derived a bound stronger than standard thermodynamic uncertainty relation (TUR) [77, 27]. In LRT, the the conventional TUR reads (in terms of the Fano factor) [57]:
@@
While, Guarnieri et al. [57] have rigorously proved that by defining a normalised probability distribution function $\widetilde{P}(\omega)$, associated with the system relaxation function $\widetilde{\Psi}_{0}(\omega)$, over the pseudomodes frequencies $\omega\in[0,\infty)$, the Fano factor can be expressed as:

### v2

@@ III.0.2 Jarzynski equality (1997):
[…]
@@
Since, for a regular protocol, the dynamical image of $D_{0}$ after time $\tau$ becomes $D_{1}$, the normalisation of $p_{f}(\Gamma_{1})$ is guaranteed. So, we can write seifert2025stochastic
@@
This remarkable property, which applies for the distribution of work and several other thermodynamic quantities, is known as the integral fluctuation theorem (IFT) udo2012stochastic ; seifert2025stochastic . From Eqs. (32)-(34), we identify
@@
with $\Delta F=F(\beta,\lambda(\tau))-F(\beta,\lambda(0))$. Substituting Eq. (35) into Eq. (34), the corresponding IFT reduces to the celebrated Jarzynski equality (JE) jarzynski1997nonequilibrium :
@@
This equality relates an average over the work required in this process to the free-energy difference between the two canonical equilibrium states: one corresponding to the value $\lambda(\tau)$ of the control parameter at time $\tau$ and other to its initial value $\lambda(0)$, both taken w.r.t. the same inverse temperature $\beta$. Further, applying Jensen’s inequality to Eq. (36), gives $\mathcal{W}=\langle W(\Gamma_{0})\rangle\geq\Delta F$ udo2012stochastic . If we define the dissipative work as $W_{\text{diss}}=W-\Delta F$, this inequality can be rewritten as $\langle W_{\text{diss}}\rangle\geq\ 0$ seifert2025stochastic . For smaller systems, unlike in the thermodynamic limit, $W_{\text{diss}}$ need not always be positive for individual realisations of the process; it may even be negative. Nevertheless, upon averaging, it becomes consistent with classical thermodynamics. If we identify $W_{\text{diss}}=T\Sigma$, where $\Sigma$ is the entropy production, then combining the IFT with Jensen’s inequality leads to $\langle\Sigma(\Gamma)\rangle\geq 0$ which is recognised as the second law of stochastic thermodynamics peliti2021stochastic . The application of JE to mesoscopic systems in modern times has been remarkable. Notably, JE places no restriction on the speed of the transformation and thus remains valid even far from equilibrium. This makes it a powerful tool for determining free-energy differences in molecular dynamics simulations, driven harmonic oscillators, and even real biomolecular experiments broedersz2022twenty . Many researchers view JE as having brought nineteenth-century thermodynamics into the twenty-first century, given its profound implications deffner2019quantumthermo .
III.0.3 Crooks Theorem (1999):
@@
$\begin{split}&\mathcal{P}_{R}(-W)=\\ =&\sum_{\Gamma^{R}_{0}\in D_{R}}\delta(W_{R}(\Gamma^{R}_{0})+W)e^{-\beta[H(\Gamma^{R}_{0},\lambda(\tau))-F(\beta,\lambda(\tau))]}\\ =&\sum_{\Gamma_{0}\in D_{F}}\delta(-W_{F}(\Gamma_{0})+W)e^{-\beta[H(\Gamma_{0},\lambda(0))+W_{F}(\Gamma_{0})-F(\beta,\lambda(\tau))]}\\ =&\mathcal{P}_{F}(W)e^{-\beta(W-\Delta F)}\end{split}$ (37)
Finally, we can write crooks1998nonequilibrium ; crooks1999entropy
@@
The CFT relation in Eq. (38) is stronger than JE [cf. Eq. (36)], since JE can be derived by integrating both sides of Eq. (38) and using $\int dW\mathcal{P}_{R}(-W)=1$. Collin et al. (2005) collin2005verification experimentally verified Crooks’ theorem in a single-molecule RNA hairpin folding–unfolding experiment using optical tweezers. In this experiment, $W_{\text{diss}}$ was sometimes negative, but on average it was positive. Thus, for small biological systems, this theorem provides a clear-cut framework for calculating free-energy changes from work distributions.
III.0.4 The different DFT and furthermore:
@@
which is an example of the detailed fluctuation theorem (DFT) seifert2025stochastic . The DFT is a stronger result than the IFT. While the IFT provides a single global constraint on the work distribution (and more generally on entropy production), the DFT establishes a local balance relation between the probabilities of observing a trajectory and its corresponding antitrajectory. Consequently, in a steady-state system, it is sufficient to know $\mathcal{P}(W)$ for positive values of $W$ in order to reconstruct the entire distribution peliti2021stochastic . Both the JE and the CFT can be regarded as special cases of the IFT and DFT when the initial and final states are equilibrium states.
@@
There exists a vast literature on FTs, depending on the observables of interest, the underlying dynamics of the system, and the hypotheses or models considered. Evans and Searles evans1994equilibrium first proposed such a relation. Subsequently, Gallavotti and Cohen gallavotti1995dynamical derived a detailed fluctuation theorem for the entropy production rate in deterministic chaotic systems. Later, the JE jarzynski1997nonequilibrium and CFT crooks1999entropy were developed for systems driven between two equilibrium states, becoming landmark results that enabled equilibrium free-energy differences to be estimated from non-equilibrium measurements. This also led to the rediscovery of earlier works by Bochkov and Kuzovlev bochkov1977general , although their definition of work differs from the present one.
@@
Moreover, introduction of trajectory probabilities for forward and backward processes naturally gave rise to a variety of FTs. For entropy production, if we separate its contributions into two parts — one arising from time dependence and another from external driving (adiabatic and non-adiabatic entropy production) — the IFT leads to the Hatano-Sasa theorem hatano2001steady , which constitutes a cornerstone in the generalisation of FDR (fluctuation-dissipation relation) to non-equilibrium steady state. Subsequently, Rao and Esposito rao2018detailed developed a unified framework that systematically derives many of the FTs reported in the literature. In addition to work and entropy, considerable attention has been devoted to the fluctuation symmetry of heat baiesi2006fluctuation ; noh2012fluctuation . Jarzynski and Wójcik jarzynski2004classical analysed the statistics of heat exchange between two systems coupled to distinct thermal baths, deriving a fluctuation relation that is now widely referred as exchange fluctuation theorem (XFT/EFT) timpanaro2019thermodynamic . More recently, the study of joint statistics of thermodynamic quantities and their associated FTs has attracted considerable attention. Interestingly, while the individual quantities may or may not obey an FT on their own, their joint probability distribution often satisfies a joint fluctuation theorem, thereby revealing richer symmetry structures in far from equilibrium thermodynamics miller2021joint ; chen2023hierarchical .
@@
So far, we have provided an account of the development of FTs in a classical settings. To understand their quantum counterparts, however, we must dig deeper. The key difference arises because any measurement of quantum states or observables disturbs the system and its trajectories, unlike in classical systems. In fact, continuous monitoring can even freeze the dynamics due to the quantum Zeno effect aulettaparisi2009quantum . In the quantum regime, the very definitions of work, heat, and entropy must be handled with care. Thus, exploring quantum versions of FTs is a rich and promising area of current research.
IV Linear Response Theory and Quantum Work Fluctuations
@@
— James Clerk Maxwell maxwell1890the
Up to this point, we have discussed the role of fluctuations in a thermodynamic setup, where general thermodynamic quantities vary due to interactions with heat bath. In 1957, Ryogo Kubo kubo1957statistical laid a concrete foundation within the Hamiltonian framework, showing how any arbitrary dynamical observable of a system coupled to a thermal bath deviates from equilibrium under an external drive, and how irreversible macroscopic behavior appears as a direct consequence of microscopic fluctuations kubo1966thefluctuation when observed through the lens of linear response of the system. Since then, Linear Response Theory (LRT) kubo1983book ; zwanzig2001nonequilibrium ; pottier2010nonequilibrium ; balakrishnan2020elements has held a central place as a powerful tool for studying near-equilibrium systems, with far-reaching applications in fluid dynamics, transport phenomena (both classical and quantum), condensed matter physics, many-body problems, chemical kinetics, and beyond nitzan2024chemical . More recently, it started elucidating its impact on quantum thermodynamics (QTD) guarnieri2024generalized ; gupt2024graph .
IV.0.1 Introduction to LRT:
Here, we establish the Kubo formulas in parallel for both classical balakrishnan2020elements ; balakrishnan2020mathematical ; kubo1983book ; pottier2010nonequilibrium ; zwanzig2001nonequilibrium and quantum mechanics kubo1983book ; pottier2010nonequilibrium . These formulas describe correlation of fluctuations, such as response and relaxation functions between observables of different dynamical variables. We discuss the properties and applications of these response and relaxation functions in the context of QTD.
@@ IV.0.1 Introduction to LRT:
We now perturb the system, starting from an initial time $t_{0}$ (eventually we take $t_{0}\to$-$\infty$) by applying a time-dependent external force $\lambda(t)$. This force is spatially homogeneous and couples with an arbitrary observable $A$ of the system. The perturbed Hamiltonian is then $H(t)=H_{0}+H^{\prime}(t)=H_{0}-\lambda(t)A$. For $t\geq t_{0}$, the density operator (or the phase-space distribution in the classical case), $\rho(t)$ evolves according to the Liouville–von Neumann equation breuer2002book ; nitzan2024chemical
@@
Here, $\{\,\cdot\,$$\cdot\,\}$ and $\cdot\,$$\cdot\,$] denote the Poisson bracket and the commutator, respectively nitzan2024chemical . This formulation allows us to treat classical and quantum dynamics in a unified manner. With this choice one recovers the standard forms:
@@
Since the Liouvillian is Hermitian (one can show $\mathcal{L}=\mathcal{L}^{\dagger}$), the operator $e^{\pm i\mathcal{L}t}$ serves as the time evolution operator kubo1983book . We now investigate the evolution of the density operator under an external perturbation in the linear regime. The Liouvillian can be split as $\mathcal{L}=\mathcal{L}_{0}+\mathcal{L}_{\text{pert}}$, associated with $H_{0}$ and $H^{\prime}$ respectively. The perturbed density operator is written as $\rho(t)=\rho_{\text{eq}}+\delta\rho(t)$. Inserting these into Eq. (43), we obtain balakrishnan2020elements ; zwanzig2001nonequilibrium
@@
where, the response function is defined as kubo1957statistical ; kubo1983book
@@
The relaxation function is then defined as kubo1957statistical ; kubo1983book
@@
which characterizes how $B$ relaxes after the perturbation is removed. Using Kubo’s identity for the canonical ensemble nitzan2024chemical ; bellac2004equilibrium ,
@@
This function is also referred to as the canonical correlation function. However, in quantum mechanics, it is often more useful to express the relaxation function in a form that accounts for the non-commutativity of operators, leading to the symmetrised correlation function, denoted as $\Psi_{AB}(t-t^{\prime})$ kubo1983book . Since it accurately captures the response (i.e., the fluctuation from equilibrium) caused by an external disturbance through time correlations between two dynamical observables, LRT has become increasingly important in quantum thermodynamics, where the control gelbwaser2015thermodynamics and optimisation of quantum devices governed by fluctuations das2024fluctuations , particularly quantum fluctuations, are of growing interest. In our context, the autocorrelation function ($A=B$), which measures how an observable correlates with itself at different times, is of central importance and will be discussed in the next section.
IV.0.2 Quanum Work Fluctuations and LRT:
@@
We live in an era where we aim to reach, utilise, and design systems on scales that have never been achieved before rossnagel2016single ; myers2022quantum ; gupt2022PRE ; sur2023quantum . To fully understand and optimise machines at the quantum level, where thermal and quantum fluctuations play equally important roles, we need to deepen our theoretical understanding of FTs, a.k.a., quantum fluctuation theorems (QFTs) hanggitalkner2015theotherqft ; funo2018quantumft ; strasberg2022quantum ; potts2024quantumthermodynamics . While QFTs can be formulated for various thermodynamic quantities, in this article we focus exclusively on work. Within the LRT framework, formulating QFTs for dissipative work has yielded new insights andrieux2008quantum . More recently, Guarnieri et al. guarnieri2024generalized demonstrated that even near equilibrium, the work distribution is strongly influenced by quantum fluctuations, which are captured through the relaxation functions.
@@
For example, consider a quantum system with Hamiltonian $H_{0}$, initially thermalised in the normalised Gibbs state $\rho_{\text{eq}}=e^{-\beta H_{0}}$. The system is then perturbed unitarily as $H_{t}=H_{0}-\lambda(t)A$ over $t\in[0,\tau]$, where $A$ is an observable and $\lambda(t)$ is a weak dimensionless driving protocol ($\lambda(t)\ll 1$, $\|A\|=1$). At $t=\tau$, when the drive is switched off, the relaxation of $A(\tau)$ is captured by the symmetrized autocorrelation function $\Psi_{0}(=\Psi_{AA})(t-t^{\prime})$, which encodes correlations of $A$ at different times. Since changes in $A$ also modify the Hamiltonian, $\Psi_{0}$ links correlations between $H_{\tau}$ and $H_{0}$, providing a natural connection to the very definition of quantum work guarnieri2024generalized ; talkner2007fluctuation .
@@
The most accomplished and widely used definition of work in QTD is the two-time projective measurement strasberg2022quantum ; deffner2019quantumthermo (TPM) scheme, where the system’s energy is measured at the beginning and end of the protocol. At $t=0$, a projective energy measurement is performed, then the system evolves under the time-dependent Hamiltonian, and at $t=\tau$ a second measurement is made. In this framework, work $W$ is a stochastic variable. For a single realization, it is given by $W[|m\rangle;|n\rangle]=E_{n}(\lambda(\tau))-E_{m}(\lambda(0))$, where $|m\rangle$ and $|n\rangle$ are the initial and final eigenstates with energies $E_{m}(\lambda(0))$ and $E_{n}(\lambda(\tau))$, respectively. The probability distribution of work is obtained by averaging over an ensemble of realizations esposito2009nonequilibrium
@@
where $p(|m\rangle\rightarrow|n\rangle)$ denotes the probability of observing the transition $|m\rangle\rightarrow|n\rangle$. If the system is initially in a thermal Gibbs state $\rho_{eq}$, Eq. (55) reduces, as shown by Scandi et al. scandi2020quantum , to the standard TPM definition of work:
$\begin{split}&\mathcal{P}(W)\\ =&\sum_{W=E_{n}-E_{m}}\langle m(\lambda(0))|\rho_{eq}|m(\lambda(0))\rangle\,\big|\langle m(\lambda(0))|n(\lambda(\tau))\rangle\big|^{2}.\end{split}$ (56)
This expression represents the product of the probability of the system initially being in state $|m\rangle$ with the transition probability $|m\rangle\rightarrow|n\rangle$. However, since each projective energy measurement perturbs the system — and may even destroy its state — the full $\mathcal{P}(W)$ cannot be directly accessed in practice through this scheme scandi2020quantum ; deffner2019quantumthermo .
However, from the statistical theory, we know that any probability distribution can be reconstructed from its cumulant generating function (CGF) strasberg2022quantum , defined as
[…]
where $\lambda=\lambda(\tau)$, denotes the final value attained by the protocol. The CGF is additive under independent random measurements, and allows one to recover the probability distribution $\mathcal{P}(W)$ through an appropriate inverse Fourier transform. Moreover, all the statistical information of $\mathcal{P}(W)$, such as cumulants, can be computed by differentiating the CGF kardar2007book :
@@
Since $W=\Delta F+W_{\text{diss}}$, the CGF naturally separates into two parts: a deterministic term depending only on the free-energy difference, and dissipative contribution due to the irreversible process, as shown by Scandi et al. scandi2020quantum
@@
[…]
Guarnieri et. al. guarnieri2024generalized , building on the elegant method of Shitara and Ueda shitara2016determining , established a landmark bridge between LRT and QFTs as
@@
where a function of $\lambda$, $g(\lambda)$, has to be convoluted with our old friend $\Psi_{0}$ to get the CGF of the dissipated work. The symmetry of $g(\lambda)=g(1-\lambda)$ ensures $K^{\text{diss}}(\lambda)=K^{\text{diss}}(1-\lambda)$ guarnieri2024generalized . Operating an inverse Laplace transform on it, gives our pre-acknowledged companion
@@
a special form of ESFT guarnieri2024generalized , now established in the quantum regime. The essential finding of Eq. (60) is that any phenomenological relaxation function can fully determine quantum work distribution in the linear regime. Moreover, this framework predicts all higher-order cumulants are positive, i.e., $\kappa_{W}^{(n)}\geq 0$ guarnieri2024generalized , leading to non-Gaussian tail (such as skewness and excess kurtosis), driven by quantum fluctuations scandi2020quantum . To quantify these non-classical effects, Guarnieri et al. guarnieri2024generalized derived a bound stronger than standard thermodynamic uncertainty relation (TUR) agarwalla2018assessing ; seifert2025stochastic . In LRT, the the conventional TUR reads (in terms of the Fano factor) guarnieri2024generalized :
@@
While, Guarnieri et al. guarnieri2024generalized have rigorously proved that by defining a normalised probability distribution function $\widetilde{P}(\omega)$, associated with the system relaxation function $\widetilde{\Psi}_{0}(\omega)$, over the pseudomodes frequencies $\omega\in[0,\infty)$, the Fano factor can be expressed as: